OpenStudy (anonymous):

step-by-Step how do i solve (x^2 + 12)/(x^2 – 2x - 35)

OpenStudy (across):

Are you trying to simplify or solve for the variable x?

OpenStudy (across):

If you're trying to simplify, I couldn't boil it down to something simpler than\[\frac{x^2+12}{x^2-2x-35}=\frac{x^2+12}{(x-7)(x+5)}\]\[\frac{x^2+12}{(x-7)(x+5)}=\frac{A}{x-7}+\frac{B}{x+5}\]\[A(x+5)+B(x-7)=x^2+12\]\[-12B=37\implies B=-\frac{37}{12}\]\[12A=61\implies A=\frac{61}{12}\]\[\frac{x^2+12}{(x-7)(x+5)}=\frac{61}{12}\frac{1}{x-7}+-\frac{37}{12}\frac{1}{x+5}\]

myininaya (myininaya):

does that work since we have an improper fraction

myininaya (myininaya):

\[\frac{x^2-2x-35}{x^2-2x-35}+\frac{2x+47}{x^2-2x-35}=1+\frac{2x+47}{x^2-2x-35}\]

myininaya (myininaya):

\[\frac{2x+47}{x^2-2x-35}=\frac{2x+47}{(x-7)(x+5)}=\frac{A}{x-7}+\frac{B}{x+5}\] \[=\frac{A(x+5)+B(x-7)}{(x-7)(x+5)}=\frac{x(A+B)+5A-7B}{(x-7)(x+5)}\]

myininaya (myininaya):

\[=>2x+47=(A+B)x+(5A-7B)\]

myininaya (myininaya):

\[A+B=2 ; 5A-7B=47\]

myininaya (myininaya):

A+B=2 5A-7B=47 7(A+B=2) 7A+7B=14 +(7A+7B=14) -------------- 12A+0=61 12A=61 A=61/12 Remember A+B=2 which implies B=2-A B=2-61/12=(24-61)/12 B=-37/12

myininaya (myininaya):

\[\frac{x^2-2x-35}{x^2-2x-35}+\frac{2x+47}{x^2-2x-35}=1+\frac{2x+47}{x^2-2x-35} \] \[=1+\frac{61}{12} \cdot \frac{1}{x-7}+\frac{-37}{12} \cdot \frac{1}{x+5}\]